OpenStudy (anonymous):

If y = 4x -2, find y when x = 4. A. 4 B. 6 C. 14

OpenStudy (anonymous):

substitute x for 4. i.e. calculate 4(4)-2.

OpenStudy (anonymous):

y= 4(4)-2 y=16-2 y=14